A sample of gas at 1.2 atm and 27 degree celcius is heated at constant pressure to 57 degree celsius lts final volume is found to be 4.75 litre. What was original volume?

Grade:12th pass

1 Answers

Arun
6 years ago
Apply the charles law
v/t of the first gas =v/t of the second
v/300=4.75/330
V = 4.75*300/330
On solving
V = 4.31 litre